SOFI data sorted to be reduced one PESSTO run at a time, with the exception of 1D reduction phase of the 2013 Oct/Nov run.
No spectroscopic SOFI targets observed during 2014 Dec/Jan, 2014 Jan/Feb, 2014 Feb/Mar, and 2014 Apr PESSTO runs.
Every night when spectroscopic science observations have been carried out with SOFI also a spectroscopic standard has been observed. However, see comment on night 20131014 below.
No problem with the wavelength calibration, RMS always ~0.2 A.
Many of the flux calibrated RG spectra of science targets and standards have noisy features in the red end of the spectrum.
Spectroscopic daytime calibrations obtained on 20130817, however, the data set does not include RG flats. SOFI spectroscopy of science targets during the 2013 Jul/Aug run only on 20130817 and with only 1 RG science target (SN2013ek). Available RG spectroscopic flats obtained during an earlier 2013 Jul/Aug subrun 20130728. However, used 20130827 RG flats from the 2013 Aug/Sep run obtained closer in time on 20130827.
Feige110 acquisition restarted twice after 1 BG exposure on 20131004 (part of an alignment test?). Used only the BG pair tFeige110_1_20131003_GB_56999_2_3.fits and tFeige110_1_20131003_GB_56999_3_2.fits in the PESSTOSOFI1dSPEC reduction phase.
No clear signal in the last exposure of the RG sequence of SN2013fc obtained on 20131014, file SN2013fc_1_20131014_GR_56999_8.fits. Unfinished cycle (9/12) reported in the night log. tSN2013fc_1_20131014_GR_56999_8_6.fits not used in the PESSTOSOFI1dSPEC reduction phase.
Feige110 RG exposures obtained on 20131014 reported to be 'probably useless due to clouds' in the night log. Visually confirmed to have a lower signal level. Opted to use both RG and BG 20131003 exposures (same run, different subrun) of Feige110 for the flux calibration of all the spectroscopic science data obtained during the 2013 Oct run, with targets observed on 20131003 and 20131014.
Available daytime calibration data set obtained on 20131209 of spectroscopic flats incomplete (only 3/4 BG flats). Spectroscopic BG flats obtained on 20131112 used to reduce the 2013 Nov/Dec run data (only 1 science target: SN2013ej BG+RG observed on 20131209).
Low signal in the last BG exposure of SN2013ek, file tSN2013ek_1_20130816_GB_56999_7_6.fits excluded from the PESSTOSOFI1dSPEC. Target observed at airmass (2.2 -> 2.9). Noisy merged spectrum SN2013ek_20130816_GB_merge_57000_1_f.fits
Cloudy night according to the night report.
No signal to trace on the SN2013fq BG+RG files. Failed to align the 2D frames (this cannot be done based on the header coordinates?), target not reduced!
SN2013fc is coincident with a background source. Selecting regions for deriving the background was not a straightforward task when extracting the spectrum. Low S/N spectra SN2013fc_20141014_GB_merge_57001_1_f.fits and SN2013fc_20131014_GR_merge_57001_1_f.fits with a strong noise feature in the red end of the RG spectrum.
SN2012ca is a faint target compared to the background. Aligning 2D frames problematic. Possibility to do this completely based on the header coordinates would be useful. Extracted spectrum dominated by background emission.
Noisy spectra SN2013fc_20131102_GR_merge_57002_1_f.fits and SN2013fc_20131102_GB_merge_57002_1_f.fits. Noise spikes also in the spectra SN2013fq_20131112_GR_merge_57003_1_f.fits
Spectroscopic standard GD71 observed twice in a row with BG and RG. Strong artifact peak in the first RG spectrum and differences in the extracted continuum shape of the two RG spectra. Used GD71_20131209_GB_merge_57002_1_f.fits and GD71_20131209_GR_merge_57002_2_f.fits to flux calibrate the extracted spectra of SN2013ej (the only spectroscopic science target observed with SOFI on 20131209).